Circular No. 5632

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

FG SAGITTAE E. F. Guinan, G. P. McCook, and T. A. Thrash, Villanova University,

report: “UBVRI photometry of FG Sge has been carried out on 86 nights since 1991 at Mt. Hopkins, AZ (0.8-m Automated Photoelectric Telescope) and at Villanova (0.38-m reflector). These observations (plus those reported on IAUC 5604, 5609, 5610, and 5619) indicate a sudden dimming of the star at all wavelengths, starting in mid-August. By early October, the star has decreased in brightness to V ≃ 12.1 and B ≃ 13.6. This is the faintest that FG Sge has been observed since the 1890s, when its photographic brightness was m_{pg} ≃ 13.1. Observations prior to the recent dimming show FG Sge to have low amplitude (∼ 0.2 mag at V), quasiperiodic (∼ 110-130 days) light variations, and a mean brightness of V ≃ 9.1-9.2. The current light loss episode could be caused by the formation of dust in an ejected shell, similar to light loss events observed in R CrB-type variables.”

GRO J0422+32 N. M. Shakhovskoj reports photometric and polarimetric data obtained

with the Crimean Astrophysical Observatory 1.25-m reflector: Aug. 26.99 UT, U = 13.46, B = 13.83, V = 13.54, R = 13.16, I = 13.04; Sept. 28.06, 13.50, 13.92, 13.57, 13.16, 12.81; Sept. 29.06, 13.46, 13.89, 13.56, 13.15, 12.76; Sept. 29.91, 13.46, 13.88, 13.54, 13.15, 12.77. Linear polarimetry in the specified bands: Aug. 16.02, U, 1.7 ± 1.6 percent, position angle 62^{o} ± 22^{o}; B, 0.1 ± 0.7, 108^{o} ± 40^{o}; V, 0.6 ± 0.9, 8^{o} ± 27^{o}; R, 0.3 ± 0.8, 13^{o} ± 36^{o}; I, 0.8 ± 0.7, 155^{o} ± 22^{o}. Aug. 26.99, U, 1.7 ± 0.5, 65^{o} ± 9^{o}; B, 0.4 ± 0.4, 77^{o} ± 23^{o}; V, 1.6 ± 0.7, 68^{o} ± 12^{o}; R, 0.7 ± 0.4, 86^{o} ± 16^{o}; I, 0.7 ± 0.5, 116^{o} ± 19^{o}. Sept. 28.07, U, 0.1 ± 0.4, 9^{o} ± 35^{o}; B, 0.5 ± 0.3, 84^{o} ± 15^{o}; V, 1.1 ± 0.5, 165^{o} ± 13^{o}; R, 0.7 ± 0.3, 87^{o} ± 10^{o}; I, 0.5 ± 0.4, 156^{o} ± 19^{o}. Sept. 29.06, U, 0.7 ± 0.3, 76^{o} ± 12^{o}; B, 0.5 ± 0.2, 90^{o} ± 10^{o}; V, 0.9 ± 0.4, 120^{o} ± 12^{o}; R, 0.7 ± 0.2, 95^{o} ± 6^{o}; I, 0.6 ± 0.3, 107^{o} ± 12^{o}. On Aug. 29, the circular polarization was measured in the R band with the 2.6-m Shajn telescope to be +0.10 ± 0.08.

SUPERNOVA 1992ba IN NGC 2082 R. H. McNaught, Anglo-Australian Observatory, reports that no image

appears on a U.K. Schmidt Telescope plate taken by A. Savage on 1992 Sept. 17.7 UT to a limit of I = 19. The comparison I survey plates were taken in 1978 and 1982.

1992 October 8 Daniel W. E. Green